OpenStudy (anonymous):

Quick Algebra 2 Problem! Medals for trying. Part 1: Is the following equation true? (2 points) 3 log[2] x+1/2 log[2] y-3 log[2] z= log[2] x^3√y/z^3 Part 2: Use complete sentences to explain the properties used in making your decision. (6 points) Long problem so therefore medals given for trying to help solve :)

Directrix (directrix):

Could you draw this? It's hard to tell which terms are where regarding numerators and denominators. Thanks

OpenStudy (anonymous):

\[3 \log_{2} x+1/2 \log_{2} y-3 \log_{2} z=\log_{2} \left(\begin{matrix}x ^{3} \sqrt{y} \\ z^{3}\end{matrix}\right)\]

OpenStudy (anonymous):

the parenthesis is obviously x^3 sqrt y over z^3 just without the line in between the two.